Identify the solution(s) of the system of equations, if any.
4x + 2y = 1
y = 5 - 3x




(–8.5, –4.5)



(–8.5, 4.5)



(–4.5, –8.5)



(4.5, –8.5)

Answer by amarjeeth123(569)

4x + 2y = 1...............equation 1
y = 5 - 3x................equation 2
Multiplying equation 2 by 2 we get,
2y=10-6x..................equation 3
Substituting equation 3 in equation 1 we get,
4x+10-6x=1
2x=9
x=4.5
Substituting in equation 2 we get,
y=5-3(4.5)=-8.5
(4.5, –8.5) is the required solution.